The device parameters for n-channel JFET are maximum drain current (IDSS) = 10mA, pinch-off voltage (Vp) =-4V

Calculate the drain current for 

(a) VGS= 0V (b) VGS=-1V (c) VGS=-4V


1
Expert's answer
2021-10-30T11:20:00-0400

The drain current of JFET is given as

(a) For VGS =0, ID =IDss = 10mA

(b) VGS= -1V, ID= 10(1-1/4)2= 3.10 mA

(c) VGS= -4V, iD = 10(1-1)2 = 0 mA